Express the roots of the equation z 3 − α 3 = 0 in terms of α and w, where w is a complex cube root of unity. Use your answer to find the roots of the following equations in the form a + ib. (a) z 3 − 27 = 0 (b) z 3 + 8 = 0 (c) z 3 − i = 0.
1
Expert's answer
2020-06-04T19:04:24-0400

Solution.

"z3\n\n\u2212\u03b13\n\n=0;"

"z^3-\alpha^3=(z-\alpha)(z^2+\alpha z+\alpha^2);"

"(z-\alpha)(z^2+\alpha z+\alpha^2)=0;"

"z−α=0 orz^2+\alpha z+\alpha^2=0;"

"z_1=\alpha;" "D=\alpha^2-4\alpha^2=3\alpha^2" ;

"z_2=\dfrac{-\alpha+\sqrt{3\alpha^2}}{2}=\alpha \dfrac{-1+i\sqrt{3}}{2};"

"z_3=\dfrac{-\alpha-\sqrt{3\alpha^2}}{2}=\alpha \dfrac{-1-i\sqrt{3}}{2};"


Cube Root of Unity Value:

"\u03c91\n\n\u200b=1 - real;"

"\omega_2= \dfrac{-1+i\sqrt{3}}{2}" - complex;

"\omega_3= \dfrac{-1-i\sqrt{3}}{2}" - complex;

"\u200bz_1=\\alpha\\sdot\\omega_1=\\alpha\\sdot1=\\alpha;"

"z_2=\alpha\sdot\omega_2=\alpha \dfrac{-1+i\sqrt{3}}{2}=-\dfrac{1}{2}\alpha+i\dfrac{\alpha\sqrt{3}}{2};"

"z_3=-\dfrac{1}{2}\alpha-i\dfrac{\alpha\sqrt{3}}{2};"

a) "z^3-27=0;"

"z_1=3;"

"z_2=-\dfrac{1}{2}\sdot3+i\dfrac{3\sqrt{3}}{2}=-\dfrac{3}{2}+i\dfrac{3\sqrt{3}}{2};"

"z_3=-\dfrac{1}{2}\sdot3-i\dfrac{3\sqrt{3}}{2}=-\dfrac{3}{2}-i\dfrac{3\sqrt{3}}{2};"

b)"z^3+8=0;"

"z_1=-2;"

"z_2=-\dfrac{1}{2}\sdot(-2)+i\dfrac{-2\sqrt{3}}{2}=-\dfrac{3}{2}-i\dfrac{2\sqrt{3}}{2};"

"z_3=-\dfrac{1}{2}\sdot(-2)-i\dfrac{-2\sqrt{3}}{2}=-\dfrac{3}{2}+i\dfrac{2\sqrt{3}}{2};"

c)"z^3-i=0;"

"z_1=i;"

"z_2=-\dfrac{1}{2}i+i\dfrac{i\sqrt{3}}{2}=-\dfrac{1}{2}i-\dfrac{\sqrt{3}}{2};"

"z_3=-\dfrac{1}{2}i-i\dfrac{i\sqrt{3}}{2}=-\dfrac{1}{2}i+\dfrac{\sqrt{3}}{2};"


Answer: a) "z_1=3;"

"z_2=-\dfrac{3}{2}+i\dfrac{3\sqrt{3}}{2};"

"z_3=-\dfrac{3}{2}-i\dfrac{3\sqrt{3}}{2};"

b)"z_1=-2;"

"z_2=-\dfrac{3}{2}-i\dfrac{2\sqrt{3}}{2};"

"z_3=-\dfrac{3}{2}+i\dfrac{2\sqrt{3}}{2};"

c)"z_1=i;"

"z_2=-\dfrac{1}{2}i-\dfrac{\sqrt{3}}{2};"

"z_3=-\dfrac{1}{2}i+\dfrac{\sqrt{3}}{2}."